A plane is travleing 480 mph with the wind, and 400 mph against the wind. Find the speed of the plane in still air.
I have no idea how to solve this problem.

Answer by mananth(16949)

plane speed in still air = x
wind speed = y
x+y=440-----------1
against wind
x-y = 400---------2
add (1) & (2)
2x=880
/2
x= 440 mph